Craig opens a savings account with $850. His account earns 3% annual simple interest. How much interest will Craig earn in 18 months?
Found 2 solutions by rfer, stanbon:
Answer by rfer(12657)

I=850*0.03*1.6
I=$40.80

Answer by stanbon(57307)

Craig opens a savings account with $850. His account earns 3% annual simple interest. How much interest will Craig earn in 18 months?
---------------------------
A(t) = P(1+nr)
-----
A(3/2) = 850(1 + (3/2)(0.03))
------
= 850 + 1.5*0.03*850
----
= $888.25
=================
Cheers,
Stan H.